OpenStudy (anonymous):

HELP?....WILL GIVE MEDAL! A house has a roof with the dimensions shown above. Both rectangular halves of the roof will be covered with plywood. If plywood comes in pieces that measure 8 feet by 4 feet, how many pieces of plywood are needed to cover the roof? (Assume that both halves of the roof are the same.)

OpenStudy (anonymous):

what are the dimensions of the roof

OpenStudy (anonymous):

@sodukogirl13

OpenStudy (anonymous):

42*52=1050 8*4=32 1050/32=32.8125 32.8125*2=65.625 65.625